OpenStudy (anonymous):

Evaluate the integral ((1/v^9)+(10/v))dv

OpenStudy (anonymous):

anti derivative of \(\frac{1}{v}\) is \(\ln(v)\) as a start

OpenStudy (anonymous):

\[\int\limits (\frac{1}{v^9}+\frac{10}{v})dv\] \[\int\limits (v^{-9}+\frac{10}{v})dv\] \[\large =-\frac{1}{8}v^{-8}+10\ln v +C\]

OpenStudy (anonymous):

\[\large =-\frac{1}{8v^8}+10\ln v +C\]

OpenStudy (anonymous):

@Azteck can you help me with (-8x+8)^-2dx I'm getting (-8x+8)^-1/8 +c ( i don't think this is wrong

OpenStudy (anonymous):

@Azteck right**

OpenStudy (anonymous):

You left out a minus sign.

OpenStudy (anonymous):

\[-\frac{1}{8}(-8x+8)^{-1}+C\] \[-\frac{1}{8(-8x+8)}+C\]

OpenStudy (anonymous):

@Azteck Thanks

OpenStudy (anonymous):

No worries.